OpenStudy (anonymous):

express using positive exponents then if possible simplify 1 over z^-6

OpenStudy (goformit100):

What is to be expressed ?

OpenStudy (anonymous):

that is what i am trying to figure out

OpenStudy (campbell_st):

it can be written as (z^{-6})^{-1} use the power of a power law \[(x^a)^b = x^{a \times b}\]

OpenStudy (anonymous):

so is that my answer I do not understand this

OpenStudy (campbell_st):

oops should be \[(z^{-6})^{-1} \]

OpenStudy (anonymous):

How do I get my answer please explain

OpenStudy (anonymous):

since the negative exponent is in the denominator, make it a positive exponent and put it in the numerator \[\frac{1}{b^{-n}}=b^n\]

OpenStudy (anonymous):

ok i got that then what?

OpenStudy (campbell_st):

use the index law for power of a power \[(x^a)^b = x^{a \times b}\] so multiply the powers you have in the question.

OpenStudy (anonymous):

ok I will try